OpenStudy (anonymous):

Simplify the following expression 6^-5

OpenStudy (anonymous):

consider x^-n=1/x^n

OpenStudy (charlotte123):

6^-5 = 1/6^5

OpenStudy (charlotte123):

Simplifying The Expression: 6^5 = 6 * 6 * 6 * 6 * 6 = 7776 So - 6^-5 = 1/6^5 = 1/7776

OpenStudy (solomonzelman):

\(\huge\color{blue}{\huge {\bbox[5pt,yellow ,border:2px solid purple]{a^{-b}=\frac{1}{a^{b}}}}}\)

OpenStudy (anonymous):

excellent

OpenStudy (solomonzelman):

\[SEE?\]

OpenStudy (charlotte123):

@SolomonZelman That box *-* CONGRATS ON 99!

OpenStudy (solomonzelman):

Oh, 99, I got it a couple hours ago...TY!

OpenStudy (charlotte123):

Late Congrats LOL And you're welcome! You deserved it!

OpenStudy (solomonzelman):

tnx. \(\huge\color{white}{\huge {\bbox[5pt,blue,border:2px solid red]{~*~~‿~*~}}}\)

OpenStudy (charlotte123):

The boxes always get me *-*

OpenStudy (solomonzelman):

\( \\ ~ \\ ~\)

OpenStudy (solomonzelman):

\[{\huge\color{(•‿•) }{\underline{I~~~~am~~~~n o t~~~~good~~~~at~~~~math~~~~ \text!}}\\ }\]

OpenStudy (charlotte123):

You are not good at Math - You are GREAT at Math THAT INVISIBLE POST THOUGH *Faints* X_X